# Two Pareto Optimum-based Heuristic Algorithms for Minimizing Tardiness and Late Jobs in the Single Machine Flowshop Problem

### v1

@@ 2.3.1 Construction of scheduling scenarios for training and testing
In order to test the heuristic insertion and train the neural network described in Section 2.2.2), it is necessary to generate a set of labeled instances for scheduling. The instances are generated with randomized arrival, execution, delay, and due times, as summarized in Table 1, which follows the notation of Section 2.1. The due times $d_{n},n=1,\ldots N$ are generated as follows:
$d_{n}=a_{n}+z_{n}+x_{n}+D^{\prime}_{n},$ (10)
@@
where the random variable $D^{\prime}_{n}$ is exponentially distributed with mean $\mu_{D^{\prime}}$. The role of $D^{\prime}_{n}$ is to allow some margin for each job, so that job $n$ can experience a net delay of $D^{\prime}_{n}$ or less and still be completed on time. The parameter values used are summarized in Table 1.
@@
Table 1: Random variables and parameters for random generation of arrival, execution, delay times, delay margin, and due times. The index $n$ labels jobs and runs from $1$ to $N$.
2.3.2 Training the neural network
@@
To improve the training performance of the NN, first the starting structure is adjusted. Preliminary investigations showed a linear activation function on the hidden layer and the stochastic gradient descent optimizer gave the best results. The batch size was not optimized, and was set at 1. The mean squared error (MSE) loss function was used to gauge performance. The number of epochs was adjusted based on the loss curves, which leads to an epoch value of $250$ where the curves begin to plateau.
@@
Figure 2: Learning loss curves for NN. The close match between training and validation curves show that overfitting has not taken place.
2.4 Performance Comparisons
In order to evaluate the relative performance of the different algorithms, a number of tests were performed. All test scenarios were generated using the distributions and parameters described in Table 1.
First we compared numerical algorithm performance to the exact MILP solution by applying MILP, NN and insertion to the same 2000 scenarios with 6 and 8 jobs, and comparing the objective functions obtained by the three methods. The number of jobs was limited because of the high runtimes of MILP.
@@ 2.4 Performance Comparisons
3 Results
3.1 Comparison of numerical algorithms with exact MILP solution for small scenarios
Figure 3 shows the distributions of objective functions obtained using the NN, MILP, and insertion heuristic for 2000 instances of 8-job scenarios with parameters given by Table 1. For the insertion heuristic since the number of jobs was limited all permutations were kept, and all insertions were included. Clearly the insertion algorithm performance is almost identical that of the exact MILP, while the NN is far less optimal.
@@ 3.1 Comparison of numerical algorithms with exact MILP solution for small scenarios
3.2 Performance comparisons between insertion and selection algorithms for larger cases
3.2.1 Objective function comparisons
@@
Figure 5 shows the difference of the objective values for several insertion algorithm variations compared to the best value for each instance. The mean and median differences are plotted, as well as the 5th and 95th percentile values. Maximum number of kept permutations varies from 30 to 80, while maximum number of insertions varies from 10 to 35. Altogether 30 combinations of parameter values are used. Results show little change in all performance measures for different kept permutation values above 50, and for insertions above 20.
Figure 6 parallels Figure 5 by showing means, medians, and 5th and 95th percentile values for objective function differences from best obtained using the selection algorithm variations. The selection algorithm with select from $7$ is far better than the selection from 6 or below. Indeed, comparison with Figure 5 shows that select from 7 also beats all insertion variants. The number of kept permutation has little effect on algorithm performance, regardless of the selection parameter.
Figure 7 and shows the proportion of times each variation of the insertion algorithm matched the overall best prediction. Each variant matches the best obtained objective value for about 30 percent of instances regardless of number of kept permutations or number of insertions. We do see a slight dip in performance for insert in the last 10 compared to larger insertion values, and an even smaller dip for number of kept permutations less than 70. This is consistent with what was seen in Figure 5, namely that the best performance is already acheived with insert in the last 15 and keeping 50 permutations.
@@
Figures 9 through 11 show pair plots that provide more detailed comparisons between algorithm variants. Figure 9 compares insertion in the last $15$, $20$, and $30$ compared to the selection with permutation windows of $5$, $6$, and $7$. The number of kept permutations is kept at $50$ for all six cases. The insertion heuristic is almost identical over the three variations, while the selection algorithm improves with a larger window. Selection with a window of $7$ performs better than the insertion variations, but insertion is better than the other two selection variations. Figure 10 compares insertion in the last $15$ and selection with a permutation window of $5$, this time over $50$, $60$, and $70$ kept permutations for each. While there is little variation over the kept permutations for either algorithm, inserting in the last $15$ is consistently better than selection with a window of $5$, especially with larger objective values. Figure 11 compares the insertion in the last $15$ and selection with a permutation window of $7$, also over $50$, $60$, and $70$ kept permutations each. While there is little variation over the kept permutations for either algorithm, inserting in the last $15$ is consistently worse than the selection algorithm with a window of $7$, especially with smaller objective values.
3.2.2 Runtime comparisons for insertion and selection
Figure 12 shows the runtime for the insertion algorithm. The runtime is linear with respect to the number of kept permutations, with more insertion slots having a steeper slope. Figure 13 shows the runtime for the selection algorithm is relatively unfazed by the number of kept permutations, where the time wavers between powers of $10$. For an extra job in the selection window, the algorithm time increases by a factor of $10$.
4 Conclusions
@@
We have shown that the performance of the insertion and selection algorithms is nearly optimal for sufficiently few jobs, and greatly improves over dispatch rules for larger numbers of jobs. For both algorithms, parameters can be adjusted to effect tradeoffs between execution time and objective function performance. In head-to-head comparisons, selection has a longer execution time when compared with a insertion algorithm with comparable accuracy. For example, The selection algorithm with a window of $6$ permutations has comparable performance with the insertion algorithm, but over 10 times the runtime. However, the insertion algorithm performance is limited and cannot be improved by further increasing the selection window and kept permutations. On the other hand, the selection algorithm can be further improved by increasing the permutation window–but at high cost in execution time.
@@
Projected future work includes more extensive benchmarking. In the literature there are other methods suggested for generating scenarios [17], which may be used to test the algorithms over a wider variety of conditions.
@@
  • Performance may perhaps be improved by combining the two algorithms. It is possible for instance that applying selection on the schedule generated by insertion may have the effect of "fine tuning", since selection performs a local reshuffling.
5 DECLARATIONS
@@
Not applicable
@@
6 Figures
@@
Figure 3: Distributions of objective function differences (NN minus MILP and insertion heuristic minus MILP) , applied to 2000 instances of 8 jobs with distributions according to Table 1.
@@
Figure 4: Cumulative distributions of objective functions for MILP, NN and insertion heuristic single-machine algorithms, applied to 2000 instances of 8 jobs with distributions according to Table 1.
@@
Figure 5: Comparison of 5th and 95th percentiles of objective minus best overall values for the insertion algorithm run with different parameter values (number of kept permutations and number of insertion slots) for 100 50-job scenarios. Best overall values were taken as the lowest objective function obtained from the 24 different insertion algorithm variants and the 18 different selection algorithm variants (see Figure 6) for the corresponding 50-job scenario. The number of insertion slots varies from 10 to 35 jobs and correspond to different series, while the number of kept permutations varies from 30 to 80 and is listed on the $x$ axis. The medians are plotted with circles and the means are plotted with $\times$’s. Insertion slots numbers 15 and and above all have similar performance, as do kept permutations 50 and above.
@@
Figure 6: Comparison of 5th to 95th percentiles of objective minus best values for the selection algorithm run with different parameter values (number of kept permutations and permutation window) for 100 50-job scenarios. Best overall values were taken as the lowest objective function obtained from the 24 different insertion algorithm variants and the 18 different selection algorithm variants (see Figure 6) for the corresponding 50-job scenario. The comparison uses selecting from 3, 4, 5, 6, and 7 job windows. Medians are plotted with circles and means are plotted with $\times$’s. Selecting from more jobs has higher accuracy. Number of kept permutations does not appear to affect the selection algorithm’s accuracy.
@@
Figure 7: Proportion that each insertion algorithm gets the minimum value out of all observed values across 50 job instances. Insertion seems to even out at around 0.3.
@@
Figure 8: Proportion that each selection algorithm gets the minimum value out of all observed values across 50 job instances. Selection continues to improve with larger windows, and reaches almost 0.5 with a selection window of 7 jobs.
Figure 9: Pair plots for insertion in the last $15$, $20$, and $30$ slots and selection with a window of $5$, $6$, and $7$, all with $50$ kept permutations. The insertion algorithms are very similar to each other, while the selection variations have more change.
Figure 10: Pair plots for insertion in the last $15$ and selection with a window of $5$, both with $50$, $60$, and $70$ kept permutations. The insertion variations are very similar to each other, and the selection variations are very similar to each other, though inserting in the last $15$ performs better than selecting with a window of $5$, especially as the objective value increases.
Figure 11: Pair plots for insertion in the last $15$ and selection with a window of $7$, both with $50$, $60$, and $70$ kept permutations. The insertion variations are very similar to each other, and the selection variations are very similar to each other, though selecting with a window of $7$ performs better than inserting in the last $15$, especially for smaller objective values.
Figure 12: Comparison of 5th to 95th percentiles of insertion algorithm runtime over 30, 40, 50, 60, 70, and 80 kept permutations using inserting in the last 10, 15, 20, 30, and 35 possible slots. The medians are plotted with circles. The time required appears linear, with a smaller slope for a smaller number of insertion slots.
Figure 13: Comparison of 5th to 95th percentiles of selection algorithm runtime over 30, 40, 50, 60, 70, and 80 kept permutations using selecting from 3, 4, 5, 6, and 7 job windows. The medians are plotted with circles. The time required appears to be mostly unaffected by the number of kept permutations. The log of runtime increases nearly linearly with the number of jobs in the selection window.

### v2

@@ 2.3.1. Construction of scheduling scenarios for training and testing
In order to test the heuristic insertion and train the neural network described in Section 2.2.2, it is necessary to generate a set of labeled instances for scheduling. The instances are generated with randomized arrival, execution, delay, and due times, as summarized in Table 1, which follows the notation of Section 2.1. The due times $d_{n},n=1,\ldots N$ are generated as follows:
$d_{n}=a_{n}+z_{n}+x_{n}+D^{\prime}_{n},$ (2.10)
@@
where the random variable $D^{\prime}_{n}$ is exponentially distributed with mean $\mu_{D^{\prime}}$. The role of $D^{\prime}_{n}$ is to allow some margin for each job, so that job $n$ can experience a net delay of $D^{\prime}_{n}$ or less and still be completed on time. The parameter values used are summarized in Table 1. All algorithms are implemented using Python 3.10, with the MILP solver from Scipy 1.12. The Spyder integrated development environment (version 5.5.1) was used.
@@
Table 1. Random variables and parameters for random generation of arrival, execution, delay times, delay margin, and due times. The index $n$ labels jobs and runs from $1$ to $N$.
2.3.2. Training the neural network
@@
To improve the training performance of the NN, first the starting structure is adjusted. In initial testing, we tried layers with between $4N$ and $8N$ nodes, and with linear, softmin, softmax, softplus, softsign, hyperbolic tangent, SELU, elu, ReLU, and sigmoid activation functions. These functions were selected because of their common use, although other activation functions exist. More limited testing on the number of hidden layers suggested that one hidden layer with $4N$ nodes and a linear activation has the best performance using the objective function listed in 2.1.2. Parameter optimization methods tried include adadelta, adagrad, adam, adamw, adamax, RMSProp, and Stochactic Gradient Descent (SGD). The batch size was not optimized, and was set at 1. The mean squared error (MSE) loss function was used to gauge performance. The number of epochs was adjusted based on the loss curves, which leads to an epoch value of $250$ where the curves begin to plateau.
@@
Figure 2. Learning loss curves for NN. The close match between training and validation curves show that overfitting has not taken place.
2.4. Performance Comparisons
In order to evaluate the relative performance of the different algorithms, a number of tests were performed. All test scenarios were generated using the distributions and parameters described in Table 1. All comparisons and training were run on a workstation with 32.0 GB of RAM and an Intel i7-1360P CPU.
First we compared numerical algorithm performance to the exact MILP solution by applying MILP, NN, and insertion to the same 2000 scenarios with 6 and 8 jobs, and comparing the objective functions obtained by the three methods. The number of jobs was limited because of the high runtimes of MILP.
@@ 2.4. Performance Comparisons
3. Results
3.1. Comparison of numerical algorithms with exact MILP solution for small scenarios
Figure 3 shows the distributions of objective functions obtained using the NN, MILP, and insertion heuristic for 2000 instances of 8-job scenarios with parameters given by Table 1. For the insertion heuristic since the number of jobs was limited all permutations were kept, and all insertions were included. Clearly the insertion algorithm performance is almost identical to that of the exact MILP, while the NN is far less optimal.
@@ 3.1. Comparison of numerical algorithms with exact MILP solution for small scenarios
3.2. Performance comparisons between insertion and selection algorithms for larger cases
3.2.1. Objective function comparisons
@@
Figure 5 shows the difference between the objective values for several insertion algorithm variants with different values of $P$ and $S$, compared to the best objective obtained for each instance over all insertion and selection algorithms. The figure shows mean and median differences for each variant, as well as the 5th and 95th percentile values. Maximum number of kept permutations varies from 30 to 80, while maximum number of insertions varies from 10 to 35. Altogether 30 combinations of parameter values are used. Results show little change in all performance measures when 50 or more permutations are kept ($P\geq 50)$, and when insertions are limited to the last 15 or more slots ($S\geq 15$). The mean values for the different algorithm variants are consistently higher than the medians, indicated by right-skewed distributions.
Figure 6 parallels Figure 5 by showing means, medians, and 5th and 95th percentile values for objective function differences from best obtained using the selection algorithm variants with different values of $K$ and $W$. The selection algorithm with $W=7$ is far better than $W\leq 6$. Indeed, comparison with Figure 5 shows that $W=7$ also beats all insertion variants. The number of kept permutations $K$ has little effect on algorithm performance, regardless of the selection parameter.
Figure 7 shows the proportion of times each insertion algorithm variant matched the overall best prediction. Each variant matches the best obtained objective value for about 30 percent of instances regardless of the values of $P$ and $S$. We do see a slight dip in performance for $S=10$ compared to $S>10$, and an even smaller dip in performane when $P<70$. This is consistent with what was seen in Figure 5, which shows that the distribution of errors for the insertion algorithm is optimized for $S\geq 10$ and $P\geq 60$.
@@
3.2.2. Pairwise comparisons of algorithm performance
Figures 9 through 11 show scatter plots that provide detailed comparisons between pairs of algorithm variants. Each point in the scatter plot represents one particular scheduling scenario . The $x$ coordinatel of the point represents the penalty achieved by one algorithm variant, and the $y$ coordinate of the point shows the penalty achieved on the same scenario by a different algorithm variant.
Figure 9 compares insertion variants with $S=15,20,30$ compared to the selection variants with $W=5,6,7$ with $P=K=50$. The 3 variations of the insertion heuristic give almost identical results (since points align closely with the 45 degree line), while the selection algorithm improves significantly as $W$ increases. Selection with a window of $7$ performs better than the insertion variations, but insertion is better than the other two selection variations.
@@ 3.2.2. Pairwise comparisons of algorithm performance
Figure 10 compares insertion variants with $S=15$ compared to the selection variants with $W=5$ with $P=50,60,70$ and $K=50,60,70$. The 3 variations of the insertion heuristic give almost identical results (since points align closely with the 45 degree line), while the selection algorithm improves significantly as $W$ increases. Selection with a window of $7$ performs better than the insertion variations, but insertion is better than the other two selection variations. While there is little variation over the kept permutations for either algorithm, inserting in the last $15$ is consistently better than selection with a window of $5$, especially with larger objective values.
Figure 11 compares insertion variants with $S=15$ compared to the selection variants with $W=7$ with $P=50,60,70$ and $K=50,60,70$. The 3 variations of the insertion heuristic give almost identical results (since points align closely with the 45 degree line), while the selection algorithm improves significantly as $W$ increases. Selection with a window of $7$ performs better than the insertion variations, but insertion is better than the other two selection variations. While there is little variation over the kept permutations for either algorithm, inserting in the last $15$ is consistently worse than the selection algorithm with a window of $7$, especially with smaller objective values.
3.2.3. Runtime comparisons for insertion and selection
Figure 12 shows runtimes for the insertion algorithm using different values for the number of kept permutations $P$ and number of insertion slots $S$. The runtime is linear $P$ for fixed $S$, with slope increasing with $S$. It appears that runtime is also linear in $S$ for fixed $P$. This linearity in parameters implies that implementations with large values of $P$ and $S$ are feasible. However, Figure 5 indicates that large values are not necessary for good performance.
Figure 13 shows the runtime for the selection algorithm with different values for permutation window $W$ and number of kept permutations $K$. The runtime is relatively constant for $K\geq 30$, wheras the runtime increases exponentially with $W$. Together with Figure 6, these results indicate that improvements in selection algorithm performance can be acheived by increasing $W$, but at relatively high computation cost.
4. Conclusions
@@
We have shown that the insertion and selection algorithms perform nearly optimally when the number of jobs is small, and offer a substantial improvement over dispatch rules as the number of jobs increases. For both algorithms, parameters can be adjusted to effect tradeoffs between execution time and objective function performance, although with insertion the performance improvements that can be achieved by parameter adjustment are limited. In head-to-head comparisons between insertion and selection, selection and insertion algorithms have comparable accuracy when comparing variants with comparable runtime. For example, The insertion algorithm with $P=50$ and $S=15$ has comparable performance and runtime with the selection algorithm with $W=6$ and $K=30$. However, the performance of the insertion algorithm is limited and cannot be improved by further increasing the selection window and kept permutations. On the other hand, the selection algorithm can be further improved by increasing the permutation window–but at high cost in execution time.
Although we did not do an extensive study of possible neural network architectures, the results from our simple architecture suggest that neural networks may not be well suited to this application. Preliminary investigations with much larger networks also showed very poor performance. It is reasonable that neural networks would have difficulty solving scheduling problems, due to their discontinuous nature. Neural networks training algorithms presuppose a continuous dependence of the loss function on inputs, so that gradient descent can be used to obtain incremental improvements. However, in scheduling problems the penalty is a discontinuous function of the inputs, so gradient-based parameter optimization cannot be expected to work.
@@
Projected future work includes more extensive benchmarking. In the literature, other methods are suggested to generate scenarios [17], which may be used to test the algorithms over a wider variety of conditions.
@@
  • Performance may perhaps be improved by combining the two algorithms. It is possible, for instance, that applying selection on the schedule generated by insertion may have the effect of ”fine tuning”, since selection performs a local reshuffling.
Finally, the two algorithms may be generalized and applied to other problems, such as the permutation flowshop problem. This research is currently in progress.
5. DECLARATIONS
@@
The code used in this paper to generate data and to apply and compare scheduling algorithms is stored in GitHub at: https://github.com/christhron/Flowshop_Iterative_Heuristics.
@@
Acknowledgements
We wish to thank the anonymous reviewers for valuable suggestions that greatly improved the clarity of the paper.
6. Figures
@@
Figure 3. Distributions of objective function differences (NN minus MILP and insertion heuristic minus MILP), applied to 2000 instances of 8 jobs with distributions according to Table 1.
@@
Figure 4. Cumulative distributions of objective functions for MILP, NN and insertion heuristic single-machine algorithms, applied to 2000 instances of 8 jobs with distributions according to Table 1.
@@
Figure 5. Performance comparison for 30 insertion algorithm variants using different parameter values (number of kept permutations $P$ and number of insertion slots $S$), for 100 50-job scenarios. The number of insertion slots ($S$) is listed on the $x$ axis, while different values of $P$ correspond to the different series shown in the legend. The $y$ axis shows the difference between the objective function obtained by the algorithm and the lowest overall objective function obtained from all 30 insertion algorithm variant shown in this figure and the 18 different selection algorithm variants shown in Figure 6 for each 50-job scenario. Error bars show the 5th and 95th percentiles for each algorithm variant. Medians for each algorithm variant are plotted with circles and are joined with lines, while means are plotted with $\times$’s. Variants with $S\geq 10$ and $P\geq 60$ have similar performance statistics.
@@
Figure 6. Comparison of 5th to 95th percentiles of objective minus best values for the selection algorithm run with different parameter values (number of kept permutations $K$ and permutation window $W$) for 100 50-job scenarios. Best overall values were taken as the lowest objective function obtained from all insertion and selection algorithm variants for the corresponding 50-job scenario. Medians are plotted with dots and connected by lines, and means are plotted with $\times$’s. Increasing $W$ produces notable reductions in error, while increasing $K$ above 30 has little effect.
@@
Figure 7. Proportion that each insertion algorithm attains the minimum penalty out of all obtained values across 50 job instances. The proportion evens out at around 0.3.
@@
Figure 8. Proportion that each selection algorithm variant attains the minimum penalty out of all obtained values across 50 job instances. Selection continues to improve with larger windows, and reaches almost 0.5 with $W=7$.
Figure 9. Scatter plots comparing penalties obtained from different insertion and selection algorithms. Pair plots for insertion algorithm variants with $S=15,20,30$ and $P=50$ and selection algorithm variants with $W=5,6,7$ and $K=50$.
Figure 10. Scatter plots comparing penalties obtained from different insertion and selection algorithms. Pair plots for insertion algorithm variants with $S=15$ and $P=50,60,70$ and selection algorithm variants with $W=5$ and $K=50,60,70$.
Figure 11. Scatter plots comparing penalties obtained from different insertion and selection algorithms. Pair plots for insertion algorithm variants with $S=15$ and $P=50,60,70$ and selection algorithm variants with $W=7$ and $K=50,60,70$.
Figure 12. Runtime comparisons for insertion algorithm variants with kept permutations $P=30,\ldots 80$ and $S=10,\ldots 35$. The median runtimes for each variant are plotted with circles, means are plotted with $\times$’s. and error bars indicate the 5th and 95th percentiles. The runtime is linear in $P$ for fixed $S$, with slope that increases as $S$ increases.
Figure 13. Runtime comparisons for selection algorithm variants with kept permutations $K=30,\ldots 80$ and permutation window size $W=5,6,7$. The median runtimes for each variant are plotted with circles, means are plotted with $\times$’s, and error bars indicate the 5th and 95th percentiles. The runtime is to be mostly unaffected by the number of kept permutations, but increases exponentially with the window size.